RC meeting and Run Plan, Friday February 17th
From SBS wiki
Return to the RC Meeting Calendar
- RC: Bogdan Wojtsekhowski, firstname.lastname@example.org
- Ran with ~30 uA beam overnight, sometimes 40 uA
- Average current was just 12 uA (8am/16 - 8am/17) => 1 C only
- CEBAF beam delivery efficiency
- Production on polarized He-3 target
- Continue running on He3 up to 45 uA if available
- Accelerator will prioritize beam and not have any maintenance days this week
Data Collection and Polarization Status
- 34.6C of uncorrected charge for this kinematic (40.6C of corrected charge)
- Last NMR 52.4%
Other Topics/Status Updates
- DAQ reboot took 1 hour during day shift
- BBcal PS multiplicity reduction?
- HCal missing channels during swing shift?
- Anything else?
- Independent alarm handler for SBS GEM system
- Bill wants to change the raster settings after some analysis
Requested Accesses to the Hall
- Jack would like to collect some cables